# Startup Acquisitions: Acquihires and Talent Hoarding

### v4

@@ 1 Introduction
This increased scrutiny is deemed unnecessary by critics arguing that startup acquisitions typically do not hamper competition – even when they result in the killing of the startup’s product or service. One common argument (e.g. Barnett, 2023) to support this view is that such acquisitions are so-called “acquihires.” As the name suggests, acquihires are essentially a hiring instrument: the acquiring firm is primarily interested in hiring the startup’s employees, not removing a potential competitor. Consider the case of Drop.io, a startup offering easy file sharing. In 2009, Drop.io was a successful startup, having been named a winner of CNET’s Webware 100 award and listed among the 50 best websites by Time magazine. After acquiring Drop.io in 2010, Facebook promptly terminated it and announced that its CEO, Sam Lessin, would be assigned to a new role. [^2] See “Webware 100 winner: Dropio,” CNET, May 2009, “50 Best Websites 2009 – Drop.io,” Time, August 2009, and “Facebook Acquires Simple File-sharing Service Drop.io,” Mashable, October 2010. While the startup was “killed,” the motivation for doing so is different from the killer acquisitions of Cunningham et al. (2021). Yet does that necessarily mean the acquisition was benign?
@@
Talent hoarding in our model is driven by the preemption effect, with firms acquiring talent partly to prevent competitors from becoming stronger. Many papers have identified the preemption effect in various settings, from patent races (Gilbert and Newbery, 1982), to M&As in the international-trade setting (Norbäck and Persson, 2004, 2007), to technology acquisitions (Bryan and Hovenkamp, 2020a), to the case of duopolists facing capacity constrained suppliers in a decentralized market (Burguet and Sákovics, 2017). Indeed, the point of our paper is to highlight that preemptive motives can be at work even when a startup operating in a different market is being acquired. However, preemptively acquiring labor is different from acquisitions of physical assets or technology because firms do not acquire property rights over labor. In our setting, this implies that some solutions to preemption, such as licensing (Katz and Shapiro, 1987), do not apply. Moreover, additional issues, such as the impact of the business cycle discussed in our third result, appear.
@@
Our final main result is that the labor-market outcomes (hiring and layoffs) for acquihired employees may become more volatile because of firms’ talent hoarding. To obtain this result, we expand our baseline model by adding a second period. In between periods, the economy may fall into a recession, and consequently, firms may get hit by potentially correlated adverse shocks. We show that relative to a benchmark case in which firms have no motive to hoard talent, talent hoarding always leads to more hiring and may also lead to more layoffs for acquihired employees when the adverse shocks are sufficiently likely or sufficiently positively correlated. This finding lends support to the view that talent hoarding was a major contributing factor to the substantial number of layoffs in the tech industry following the increased downward pressure on the US economy in 2022.333See “Tech’s Talent Wars Have Come Back to Bite It,” New York Times, November 2022.
Talent hoarding is likely to be a problem whenever firms can identify highly impactful individuals before their competitors do so. As such, it can, in principle, also materialize when firms hire new employees who possess a scarce and desirable skill set. In such situations, the insights of our model will also apply. However, we focus on acquihires (and not hiring in general) because talent hoarding is likely to be particularly acute in this setting, as it combines individuals who may not be known in the industry yet (the entrepreneurs) with a verifiable signal of their ability (the startup that they built).
@@
We explore two important extensions in Section 5. First, we consider the situation where a startup holds value for firms due to both its employees and its technology. This extension highlights that, unlike labor, technology can be sold or licensed following an acquisition since firms do not obtain property rights over labor. Second, we examine partial acquisitions or investments in startups. In these instances, the acquiring firm purchases only a portion of the startup, gaining a share of its profits and some decision-making authority. We demonstrate that when this decision power includes the ability to block acquisitions by competitors, such investments can be an attractive alternative to full acquihires. We also show that partial acquisitions generate lower inefficiencies than full acquihires but that they occur more often. Beyond these two extensions, we discuss in Section 6 how the emergence of talent hoarding is robust to incorporating asymmetry in market power among potential acquirers, expanding the number of firms, introducing uncertainty in the order of moves, and allowing for some surplus sharing between the acquiring firm and the entrepreneur. Overall, these extensions not only confirm the robustness of our core findings but provide additional insights. For instance, our extension with asymmetric firms highlights that, under reasonable profit assumptions, dominant firms have stronger incentives to hoard talent – an observation that resonates with concerns expressed by many regulators regarding the excessive hiring prevalent among tech giants.
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2020, 2023; Katz, 2021; Letina et al., 2024). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2024), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or heavily investing in innovation (Kamepalli et al., 2021; Teh et al., 2022; Shelegia and Motta, 2021; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2023; Eisfeld, 2022; Gautier and Lamesch, 2021). Our paper differs from this literature by considering startups that are not potential competitors of the incumbents. The main channel through which acquisitions create inefficiencies is thus fundamentally different.
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2023) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly out of a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2023; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 2 Model
From the consumers’ point of view, there are three possible outcomes. First, absent an acquihire all three firms are active in their respective markets. In this case, the consumer surplus arising from the competition between the two symmetric firms is $CS_{F}$ and that from the startup is $CS_{E}$. Second, a low-match acquihire results in competition between the two (now asymmetric) firms generating the entire consumer surplus ($CS_{L}$). Third, a high-match acquihire results in a similar competition that generates the entire consumer surplus ($CS_{H}$). Whenever an acquihire occurs, the consumer surplus generated by the startup ($CS_{E}$) is lost. The next assumption captures the idea that as one of the two firms becomes more efficient, it passes on some of that efficiency gain to consumers. [^5] While it is possible that this assumption does not hold, extending our analysis to those cases is straightforward but would come at the cost of more complex exposition.
@@
We expect our model to be more applicable to situations in which the startup is still relatively young and not widely known, so that firms becoming aware of it sequentially is appropriate. Indeed, as evidenced in the report issued by the Federal Trade Commission’s report on non-Hart-Scott-Rodino acquisitions by Alphabet, Amazon.com, Apple, (then still) Facebook, and Microsoft from 2010-2019, such situations seem to be quite prevalent if not the norm (FTC, 2021). The report shows that (i) approximately 65% of all transactions cost less than \$25m, (ii) almost half of all transactions comprised startups that were less than five years old, (iii) more than 50% of the transactions involved targets with fewer than ten employees. Overall, this supports the notion of small, relatively young and not widely known startups. [^7] Should the startup already be more mature or the founder(s) actively looking for an exit, a simultaneous-move model may be more suitable. In such instances, talent hoarding, as outlined in Proposition 1 below, is less likely to occur, as a firm with high match value should be able to outbid a low-match-value competitor.
@@ 3 Talent Hoarding
The result in Proposition 1 shows that talent hoarding may occur when (i) $\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}>\pi_{E}$ so that $\lambda_{A}<1$, and (ii) the probability of a high match is sufficiently high. The condition $\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}>\pi_{E}$ guarantees that the gain for a low-match firm from an acquihire when facing a high-match competitor is bigger than the cost of the acquihire. However, since a low-match firm makes a negative profit from the acquihire per se, it will only proceed when facing a high-match competitor is likely enough. [^8] In classical labor models where firm-employee match value matters (e.g., Jovanovic, 1979), firms only care about their own match value. In our model, because of oligopolistic competition, the potential match value between the competitor and the worker is driving the results. Effectively, a low-match firm 1 is willing to overpay when making the acquihire to prevent the potential emergence of a highly competitive firm 2. While a low-match firm 1 does reap some efficiency gain from the acquihire, it is the threat of a more competitive firm 2 that motivates the acquihire. Thus, talent hoarding is more likely if the price of the acquisition $\pi_{E}$ is low and the probability of a high-match competitor $\lambda$ is high. [^9] One may wonder why the low-match firm does not keep the startup operational or what would happen if the startup was also valuable because of its technology. We discuss these issues in Section 5. Briefly, we argue that operating the startup as a subsidiary might be less profitable than integration due to moral-hazard issues. We also show that when startups own valuable technology, the firms have no incentive to hoard technology but still hoard talent.

### v5

@@ 1 Introduction
This increased scrutiny is deemed unnecessary by critics arguing that startup acquisitions typically do not hamper competition – even when they result in the killing of the startup’s product or service. One common argument (e.g. Barnett, 2023) to support this view is that such acquisitions are so-called “acquihires.” As the name suggests, acquihires are essentially a hiring instrument: the acquiring firm is primarily interested in hiring the startup’s employees, not removing a potential competitor. Consider the case of Drop.io, a startup offering easy file sharing. In 2009, Drop.io was a successful startup, having been named a winner of CNET’s Webware 100 award and listed among the 50 best websites by Time magazine. After acquiring Drop.io in 2010, Facebook promptly terminated it and announced that its CEO, Sam Lessin, would be assigned to a new role (Webware staff, 2009; Fisher, 2009; McCarthy, 2010). While the startup was “killed,” the motivation for doing so is different from the killer acquisitions of Cunningham et al. (2021). Yet, does that necessarily mean the acquisition was benign?
@@
Talent hoarding in our model is driven by the preemption effect, with firms acquiring talent partly to prevent competitors from becoming stronger. Many papers have identified the preemption effect in various settings, from patent races (Gilbert and Newbery, 1982), to M&As in the international-trade setting (Norbäck and Persson, 2004, 2007), to technology acquisitions (Bryan and Hovenkamp, 2020a), to the case of duopolists facing capacity constrained suppliers in a decentralized market (Burguet and Sákovics, 2017). Indeed, the point of our paper is to highlight that preemptive motives can be at work even when a startup operating in a different market is being acquired.
However, preemptively acquiring labor is different – and can be more pernicious – from acquisitions of physical assets or technology, because firms do not acquire property rights over labor. In particular, when a firm preemptively obtains physical assets or technology that could be more profitably used by the competitor, both firms have an incentive to negotiate an efficient allocation of property rights, for example by licensing the technology, as in Katz and Shapiro (1987). This is not the case when a firm preemptively obtains labor. In this case, the best that the firm can do is to try to keep the workers from joining the competitor, thereby leading to an inefficient allocation of resources. A similar misallocation of labor can also occur within a firm (and therefore without a competition motive), as Haegele (2022) has documented. Moreover, additional issues, such as the impact of the business cycle, which we will discuss later, appear.
@@
An important question is the extent to which the emergence of talent hoarding and its welfare implications hinge on our modeling assumptions. To this end, Section 4 considers several variations of our model including changes to the order and timing of moves, bargaining power, or the market’s competitiveness. The key takeaway is that talent hoarding is not an artifact of our model but that the preemption motives and their adverse effects are a robust phenomenon. However, the section’s results also highlight that there are situations in which inefficient acquihires are less of a concern and do not necessarily warrant regulatory attention.
@@
Our final main result is that the labor-market outcomes (hiring and layoffs) for acquihired employees may become more volatile because of firms’ talent hoarding. To obtain this result, we expand our baseline model by adding a second period. In between periods, the economy may fall into a recession and, consequently, firms may get hit by potentially correlated adverse shocks. We show that relative to a benchmark case in which firms have no motive to hoard talent, talent hoarding always leads to more hiring and may also lead to more layoffs for acquihired employees when the adverse shocks are sufficiently likely or sufficiently positively correlated. This finding lends support to the view that talent hoarding was a contributing factor to the substantial number of layoffs in the tech industry following the increased downward pressure on the US economy in 2022 (Griffith, 2022).
@@
We extend our framework in two important directions in Section 6. First, we consider the situation where a startup holds value for firms due to both its employees and its technology. This extension highlights the insight that, unlike labor, technology can be sold or licensed following an acquisition since firms do not obtain property rights over labor. Second, we examine partial acquisitions or investments in startups. In these instances, the acquiring firm purchases only a portion of the startup, gaining a share of its profits and some decision-making authority. We demonstrate that when this decision power includes the ability to block acquisitions by competitors, such investments can be an attractive alternative to full acquihires. We also show that partial acquisitions generate lower inefficiencies than full acquihires but that they occur more often.
We close the introduction by situating acquihires within the broader context of labor market dynamics and hiring practices. Talent hoarding, as described in our model, is most likely to arise when firms can identify high-impact teams before their competitors. This mechanism could, in principle, extend beyond acquihires to other forms of hiring, such as direct recruitment of employees with scarce and desirable skills. However, we focus on acquihires but not hiring in general because they present a particularly acute setting for talent hoarding: startups provide a clear and verifiable signal of entrepreneurial ability, while the individuals behind them may not yet be widely recognized within the industry. That said, firms may also choose to bypass the acquisition process altogether by directly poaching a startup’s employees (see Bar-Isaac et al., 2023, and the pertaining discussion in the literature review). Indeed, this practice has recently been documented in the context of AI startups and been portrayed as a way to avoid regulator scrutiny that would ensue with an acquisition of the startup (Deutscher, 2024; Heath, 2024). For example, in the case of Microsoft poaching employees from the AI startup Inflection, the Competition and Markets Authority in the UK investigated the transaction and treated it as a merger. Such hiring events can have similar economic and competitive implications to acquihires. Indeed, as the Microsoft/Inflection case suggests, these events may affect the acquirer, the hired employees, and the startup in ways comparable to an acquihire, warranting equivalent regulatory treatment (Madhumita and Hodgson, 2024).
@@
Our paper is most closely related to the literature studying the economics of startup acquisitions. Much of the early literature examined, in various settings, how the prospect of an acquisition affects the incentives of startups and incumbents to invest in innovation (e.g., Mason and Weeds, 2013; Norbäck and Persson, 2012; Gans and Stern, 2000; Phillips and Zhdanov, 2013; Rasmusen, 1988). Following Cunningham et al. (2021), who demonstrated that incumbents may acquire startups for anticompetitive reasons, a large literature has studied the effects a more restrictive merger policy would have on innovation and overall welfare (Motta and Peitz, 2021; Cabral, 2020, 2023; Katz, 2021; Letina et al., 2024). Others examine how acquisitions can steer the direction of innovation (Bryan and Hovenkamp, 2020a; Callander and Matouschek, 2022; Dijk et al., 2024), and Fumagalli et al. (2023) consider the impact of financial constraints. Several papers consider dynamic incentives (Cabral, 2018; Bryan and Hovenkamp, 2020a; Hollenbeck, 2020; Denicolo and Polo, 2021). A key insight is that if the incumbent pulls too far ahead in the technology space, the pace of innovation will go down. There is also the possibility that the incumbent creates a kill-zone that disincentivizes entry, either by acquiring entrants, copying their products, or investing heavily in innovation (Kamepalli et al., 2021; Teh et al., 2022; Shelegia and Motta, 2021; Bao and Eeckhout, 2023). On the empirical side, Ederer and Pellegrino (2023) show that startups increasingly favor acquisitions over IPOs as exit strategies. Finally, several papers empirically study acquisitions in the tech sector (Affeldt and Kesler, 2021a, b; Prado and Bauer, 2022; Jin et al., 2023; Gugler et al., 2023; Eisfeld, 2022; Gautier and Lamesch, 2021).
Our paper differs from this literature by considering startups that are not potential competitors of the incumbents. The main channel through which acquisitions create inefficiencies is thus fundamentally different. In particular, the welfare loss does not occur due to the loss of competition, but rather due to inefficient allocation of talent. For regulators, this implies that it is not sufficient to show that there is no (potential) product market overlap between the target and the acquirer for the acquisition to be benign – as it would be if only the concerns from the literature on potential competitors were present. Finally, while in the potential competitor model the anticompetitve concerns are typically strongest when there is a single incumbent, in our model this is not the case. At least two incumbents are needed in our model for the inefficiency to arise, and, as we show later on, the incentives to hoard talent can initially even increase as the number of incumbents grows beyond two.
@@
Our paper also relates to the empirical literature that directly studies acquihires. Ouimet and Zarutskie (2020), Ng and Stuart (2021), Chen et al. (2021), and Chen et al. (2023) show that acquiring talent is indeed an important motivation for acquisitions. However, acquihired employees separate at a higher rate than regularly hired employees (Ng and Stuart, 2021; Verginer et al., 2022), possibly due to a preference for working at startups or a misalignment with the acquirer’s plans (Kim, 2023; Loh et al., 2019). This empirical finding is consistent with our theoretical result.
@@ 2 Model
From the consumers’ point of view, there are three possible outcomes. First, absent an acquihire all three firms are active in their respective markets. In this case, the consumer surplus arising from the competition between the two symmetric firms is $CS_{F}$ and that from the startup is $CS_{E}$. Second, a low-match acquihire results in competition between the two (now asymmetric) firms generating the entire consumer surplus ($CS_{L}$). Third, a high-match acquihire results in a similar competition that generates the entire consumer surplus ($CS_{H}$). Whenever an acquihire occurs, the consumer surplus generated by the startup ($CS_{E}$) in its original market is lost, although the market itself need not disappear entirely. [^3] The reduction in consumer surplus due to the startup’s disappearance can be substantial, even if the startup’s profits are not. For an example of such a “loved but unprofitable” startup, consider the mapping app Zenly, which had more than 15 million daily users when it was controversially shut down by its parent company based on the “low likelihood of the app ever turning a meaningful profit” (O’Brien, 2022). The next assumption captures the idea that as one of the two firms becomes more efficient, it passes on some of that efficiency gain to consumers. [^4] While it is possible that this assumption does not hold, extending our analysis to those cases is straightforward but would come at the cost of more complex exposition.
@@
We expect our model to be more applicable to situations in which the startup is still relatively young and not widely recognized, making the assumption that firms become aware of it sequentially more realistic. Indeed, as evidenced in the Federal Trade Commission’s report on non-Hart-Scott-Rodino acquisitions by Alphabet, Amazon.com, Apple, (then still) Facebook, and Microsoft from 2010-2019, such situations seem to be quite prevalent if not the norm (FTC, 2021). The report shows that (i) approximately 65% of all transactions cost less than \$25m, (ii) almost half of all transactions comprised startups that were less than five years old, (iii) more than 50% of the transactions involved targets with fewer than ten employees. Overall, this supports the notion of small, relatively young and not widely known startups. [^6] Should the startup already be more mature or the founder(s) actively looking for an exit, a simultaneous-move model might be more appropriate. We discuss this alternative setting and its implications on the prevalence of talent hoarding in Section 4. Moreover, while the match quality $\theta$ is the acquirer’s private information, we view this as information that has to be learned by interacting with and getting to know the startup. Consequently, a firm that does not yet know the startup (well enough) cannot engage in costly signaling to prevent a low-type competitor of preemptively acquiring the startup. [^7] Even if a high-type acquirer could signal its type, a classical hold-up problem may arise, as it is unclear that it could credibly commit to paying more than a low-type acquirer. Indeed, it is a priori uncertain which of the two types would have a higher willingness to pay for the startup, as can be seen from the discussion of a simultaneous-move variant in Section 4.
@@ 3 Talent Hoarding
The result in Proposition 1 shows that talent hoarding may occur when (i) $\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}>\pi_{E}$ so that $\lambda_{A}<1$, and (ii) the probability of a high match is sufficiently high. The condition $\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}>\pi_{E}$ guarantees that the gain for a low-match firm from an acquihire when facing a high-match competitor is bigger than the cost of the acquihire. However, since a low-match firm makes a negative profit from the acquihire per se, it will only proceed when facing a high-match competitor is likely enough. [^8] In classical labor models where firm-employee match value matters (e.g., Jovanovic, 1979), firms only care about their own match value. In our model, because of oligopolistic competition, the potential match value between the competitor and the worker is driving the results. Effectively, a low-match firm 1 is willing to overpay when making the acquihire to prevent the potential emergence of a highly competitive firm 2. While a low-match firm 1 does reap some efficiency gain from the acquihire, it is the threat of a more competitive firm 2 that motivates the acquihire. Thus, talent hoarding is more likely if the price of the acquisition $\pi_{E}$ is low and the probability of a high-match competitor $\lambda$ is high. [^9] One may wonder why the low-match firm does not keep the startup operational or what would happen if the startup was also valuable because of its technology. We discuss these issues in Section 6. Briefly, we argue that operating the startup as a subsidiary might be less profitable than integration due to moral-hazard issues. Indeed, after the acquisition of Zenly by Snap (see footnote 3), the startup was kept essentially independent. One can interpret the owner’s decision to shut down the startup as recognizing that Zenly’s autonomy encouraged risky expansion over profitability. Further, we also show that when startups own valuable technology, the firms have no incentive to hoard technology but still hoard talent.
@@
4 Robustness and Limits to Regulation
Propositions 1 and 2 provide conditions – within our baseline framework – under which talent hoarding is likely to occur and when acquihires are harmful for consumer surplus, respectively. Taken together, these results also shed light on the welfare impact of talent hoarding. However, it is essential to assess the extent to which these findings hold beyond our baseline framework. To this end, we explore several variations of our benchmark model. Through these extensions, we clarify the robustness and limitations of our results regarding the prevalence of talent hoarding and discuss additional factors that policymakers should consider when regulating acquihires. The formal analysis underlying this section is presented in Appendix B.
Simultaneous Moves.
When both firms are aware of the startup, as would be the case when the entrepreneur is already established and well-known, modeling a simultaneous-move game might be more appropriate. This scenario effectively corresponds to an auction with externalities, as described by Jehiel et al. (1996). Although not all of their results are directly applicable here, some insights are relevant. To begin with, consider the case of complete information and assume that one firm has a high and the other a low type. Then, the entrepreneur optimally sells the startup to the low-type firm if and only if
$\bar{\Pi}_{F}^{L}+\underline{\Pi}_{F}^{L}\geq\bar{\Pi}_{F}^{H}+\underline{\Pi}_{F}^{H}.$ (3)
Hence, the low-type firm outbids the high-type firm whenever the sum of their profits is greater after a low-type acquisition than after a high-type acquisition. [^11] This condition is unlikely to hold in standard models with homogeneous goods, and in particular it is not satisfied in our Cournot example. However, it can hold in models with differentiated goods where each firm has a large base of “loyal” consumers. While this outcome is efficient from an industry-wide perspective, talent hoarding still arises as the workers would be more efficiently employed at the startup. Conversely, when the inequality in (3) is reversed, the high-type firm outbids the low-type firm, thereby eliminating talent hoarding. Extending this analysis to an incomplete-information setting does not alter the qualitative insights.
Overall, in simultaneous-move settings, talent hoarding remains a possibility but is likely to be less prevalent than in our benchmark model. Consequently, the associated welfare concerns are attenuated, suggesting that talent hoarding may not be a strong justification for blocking startup acquisitions in these situations.
Uncertain Order of Moves.
As previously noted, talent hoarding is more likely to occur when firms can identify high-impact individuals before their competitors. However, when a firm becomes aware of a startup with such employees, it may not know whether its competitors are also aware. To capture this uncertainty, we extend the model by introducing an unknown order of moves, where the sequence is privately determined by nature with uniform probabilities. In this setting, a symmetric equilibrium exists in which all firms undertake an acquihire regardless of their type if
$\lambda\geq\lambda_{A,U}\equiv\frac{\pi_{E}+\underline{\Pi}^{L}_{F}-\bar{\Pi}_{F}^{L}}{\underline{\Pi}^{L}_{F}-\underline{\Pi}_{F}^{H}}.$ (4)
This condition is more stringent than in the benchmark model, but talent hoarding may still arise. Nonetheless, the likelihood of adverse welfare effects diminishes, as the uncertainty surrounding the order of moves reduces the prevalence of talent hoarding.
Emerging Competitor.
We now consider a scenario where the two incumbents do not co-exist. That is, only one incumbent is initially active, but the other firm may yet enter. [^12] This setting with a potentially emerging competitor is reminiscent of Segal and Whinston (2000). However, allowing for ex-post renegotiation with an emerging competitor, as in their paper, has an ambiguous effect on the incentives to hoard talent. If a low-type firm could sell the startup to an emerging high-type firm, the incentives to acquire the startup may increase, as the loss incurred through the acquisition can be recouped when selling to the high-type competitor. At the same time, the possibility that the competitor does not even emerge reduces the incentives to acquihire the startup in the first place, leaving the total effect ambiguous. Specifically, the entry of the second firm only happens with probability $q\in[0,1]$. If the second firm does not enter, and there is no acquisition, the first firm makes a monopoly profit $\Pi_{M}>\Pi_{F}$. Acquihiring the startup yields a match-dependent monopoly profit of $\bar{\Pi}_{M}^{H}$ or $\bar{\Pi}_{M}^{L}$, where—mirroring Assumption 1—we assume $\bar{\Pi}^{H}_{M}>\Pi_{M}+\pi_{E}>\bar{\Pi}^{L}_{M}$. If the second firm does enter, the same profits as in the benchmark accrue (with or without an acquihire). In the case of an emerging competitor, the condition for talent hoarding to emerge is given by
$\lambda\geq\lambda_{A,M}\equiv\frac{\pi_{E}-q(\bar{\Pi}_{F}^{L}-\Pi_{F})-(1-q)(\bar{\Pi}_{M}^{L}-\Pi_{M})}{q(\Pi_{F}-\underline{\Pi}_{F}^{H})}.$ (5)
As $q$ increases, this condition approaches condition (1) from the benchmark model. However, as $q$ decreases – and thus the threat of an entrant who might acquihire the startup diminishes – the condition becomes less likely to hold. In such a situation, adverse welfare effects are less likely than in the benchmark, as talent hoarding is less likely to occur and consumer surplus is less likely to be reduced. Finally, as the entry of the competitor is uncertain to begin with, the regulator ought to be cautious when considering the ban of startup acquisitions.
Surplus Sharing
In our baseline model, the acquiring firm has all the bargaining power and thus manages to acquire the startup at a price that leaves the entrepreneur with zero surplus. In this extension, we show that talent hoarding can persist even when surplus is shared, provided the entrepreneur’s bargaining power is not too large. Formally, we assume that in case of an acquihire the entrepreneur receives a share $\xi\in[0,1]$ of the surplus. Then, the condition for talent hoarding becomes
$\lambda\geq\lambda_{A,S}\equiv\frac{\pi_{E}+\Pi_{F}-\bar{\Pi}_{F}^{L}}{\Pi_{F}-\underline{\Pi}_{F}^{H}-\xi(\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E})}.$ (6)
Evidently, as $\xi$ and thus the entrepreneur’s bargaining power increases, the condition is less likely to be satisfied. In particular, if
$\xi>\frac{\bar{\Pi}_{F}^{L}-\underline{\Pi}_{F}^{H}-\pi_{E}}{\bar{\Pi}_{F}^{H}-\Pi_{F}-\pi_{E}}.$
the acquisition of the startup becomes prohibitively expensive for the firm and talent hoarding does not emerge in equilibrium. Once again, adverse welfare impacts are less likely. The result suggests that regulators could try to use estimates of the bargaining power in startup acquisitions to gauge whether an acquisition should be blocked or not.
Dominant Firm.
Instead of symmetric firms, the market could be characterized by a dominant firm and a challenger. We show that the incentives to hoard talent also emerge in this asymmetric setting. Specifically, suppose that the dominant firm both stands to lose more if a startup is acquired by the challenger (maybe because the dominant firm has a larger market share, so it has more to lose) and stands to gain more by acquihiring the startup itself (e.g., because its larger market share allows it to deploy improvements more rapidly to a broader consumer base). Moreover, assuming that a dominant firm makes more profit than symmetric firms (i.e., $\Pi_{D}\geq\Pi_{F}$) and is more likely to identify startups than its competitor, then talent hoarding is more likely to emerge than in the benchmark model, as the relevant condition now reads
$\lambda>\lambda_{A,D}\equiv\frac{\pi_{E}+\Pi_{D}-\bar{\Pi}_{D}^{L}}{\Pi_{D}-\underline{\Pi}_{D}^{H}},$ (7)
which is less stringent than condition (1). This suggests that acquisitions by dominant firms may deserve more scrutiny by the regulator, as they are more likely to lead to talent hoarding and negative effects on consumer welfare. However, an outright ban on such acquisitions may not be justified, as acquisitions by challenger firms are less likely to produce these adverse outcomes.
Multiple Firms.
Finally, we examine markets with multiple firms that may sequentially attempt to acquihire the startup. Focusing on the Cournot-Oligopoly setting, we show that in the limit, as the number of firms grows large, talent hoarding ceases to take place. Intuitively, the profit at risk from a competitor’s acquihire becomes smaller as the number of competitors increases, reducing the incentive to engage in costly talent hoarding. However, the effect of the increase in the number of competitors on the incentive to hoard talent is not necessarily monotonic. The reason for the non-monotonicity is that an increase in the number of competitors, in addition to decreasing profit at risk, also increases the probability that a high-match competitor will acquire the startup. For a small number of firms, it is unclear which effect dominates. Indeed, we show in a parametric example that talent hoarding may be more prevalent with three firms than with two. Even if talent hoarding may arise with more than two firms, the concerns about adverse welfare effects should be less pronounced when markets are more competitive.